OpenStudy (anonymous):

Rewrite with rational exponents. (√(3mn))^7

jimthompson5910 (jim_thompson5910):

\[\Large \left(\sqrt{3mn}\right)^7\] \[\Large \left((3mn)^{\frac{1}{2}}\right)^7\] \[\Large (3mn)^{\frac{7}{2}}\] So \[\Large \left(\sqrt{3mn}\right)^7 = (3mn)^{\frac{7}{2}}\]